# Neural Diffusion Processes for Physically Interpretable Survival Prediction

### v2

@@ V.2 Physics-based interpretability
which assigns to each point $\mathbf{p}$ in the parameter space an estimated event time $T(\mathbf{p})$. These values serve as proxies for patients’ risk and allow us to test whether the model implements a similarity notion based on distance in parameter space.
@@
Figure 4: Event times in the parameter spaces of Deep FHT models. Left: FLChain dataset in the space of Lévy Deep FHT ($\{x_{0},D\}$). Right: Framingham dataset in the space of inverse Gaussian Deep FHT ($\{x_{0},\mu\}$). Background colors represent interpolated event times $T(\mathbf{p})$ obtained by inverse distance weighting of uncensored training instances. Points correspond to uncensored test patients and individuals surviving beyond the last uncensored time, colored by their observed event times. In both cases, patients with similar event times cluster in contiguous regions, illustrating that the model encodes similarity in terms of process parameters.
@@
In Fig. 4, a colormap based on $T(\mathbf{p})$ is superimposed on the parameter space, with uncensored test instances and individuals surviving beyond the last uncensored time represented as points colored by their observed event times. The plots show that previously unseen patients are typically assigned to regions with comparable event times, indicating that individuals with similar risk cluster together in the space of process parameters.
@@
We also note that the empirical distribution of event times shown in Fig. 4 — and used as a proxy for patient risk — matches our theoretical knowledge of process dynamics. Patients with longer survival times cluster in regions with favorable initial conditions (i.e., larger distance from the absorbing barrier) and lower diffusion coefficient or drift; processes with these parameter values typically yield longer first-passage times. Conversely, patients experiencing early failure are assigned higher $D$ or $\mu$ and initial conditions closer to the barrier, consistent with shorter first-passage times.
@@
Final remarks.  We conclude this section by noting a phenomenon visible in Figs. 4,  5: the apparent collinearity between process parameters (or their logarithms). This effect, observed in both Lévy and inverse Gaussian models, leads to data points concentrating along one-dimensional manifolds in parameter space. As reported in the FHT literature [12], this behavior can be attributed to the limited complexity of publicly available datasets rather than to intrinsic shortcomings of the models themselves. We reinforce this notion by showing that the parameter space of the synthetic NonPH dataset does not exhibit this effect (see Appendix B, Fig. 6).
Figure 5: Feature–parameter relationships in the Lévy and inverse Gaussian DeepFHT models. Top: Framingham dataset with Lévy model, showing the parameter space $\{x_{0},D\}$ colored by systolic (left) and diastolic (right) blood pressure. Bottom: GBSG2 dataset with inverse Gaussian model, showing the parameter space $\{x_{0},\mu\}$ for patients with grade 1 (left) and grade 3 (right) tumors.
In both cases, clinical risk factors align with model-derived high-risk regions (small $x_{0}$, large $D$ or $\mu$), supporting the physical interpretability of the parameterization.
@@ Appendix A Model architecture and hyperparameters
Model GBSG2 Framingham SUPPORT2 FLChain NonPH
Lévy $0.694\pm 0.017$ $0.705\pm 0.012$ $0.854\pm 0.004$ $0.929\pm 0.003$ $0.840\pm 0.004$
Inverse Gaussian $0.692\pm 0.024$ $0.700\pm 0.013$ $0.854\pm 0.005$ $0.926\pm 0.003$ $0.837\pm 0.010$
@@
Appendix B NonPH parameter space
Figure 6: NonPH test set in the parameter spaces of Lévy and inverse Gaussian DeepFHT models. Notice the absence of collinearity between parameters, as discussed in Sec. V.2
Appendix C Synthetic dataset generation

### v3

@@ V.2 Empirical analysis of parameter-space interpretability
Event times in parameter space.  Using the IDW interpolation defined in Section IV.1, we map each point in the parameter spaces to an estimated event time $T(\mathbf{p})$, computed from uncensored training instances. These maps serve as risk proxies and allow to test whether the model encodes similarity as proximity in parameter space.
@@
Figure 3: Event times in the parameter spaces of Deep FHT models. Left: Framingham dataset in the space of Lévy Deep FHT ($\{x_{0},D\}$). Right: Framingham dataset in the space of inverse Gaussian Deep FHT ($\{x_{0},\mu\}$). Background colors represent interpolated event times $T(\mathbf{p})$ obtained by inverse distance weighting of uncensored training instances. Points correspond to uncensored test patients and individuals surviving beyond the last uncensored time, colored by their observed event times. In both cases, patients with similar event times cluster in contiguous regions, illustrating that the model encodes similarity in terms of process parameters.
@@
In Fig. 3, a colormap based on $T(\mathbf{p})$ is superimposed on the parameter space of both Lévy and inverse Gaussian models trained on the Framingham dataset, with uncensored test instances and individuals surviving beyond the last uncensored time represented as points colored by their observed event times. The plots show that previously unseen patients are typically assigned to regions with comparable event times, indicating that individuals with similar risk cluster together in the space of process parameters. This proves to be true across all datasets and models (see Appendix C).
@@
We also note that the empirical distribution of event times shown in Fig. 3 matches our theoretical knowledge of process dynamics. Patients with longer survival times cluster in regions with favorable initial conditions (i.e., larger distance from the absorbing barrier) and lower diffusion coefficient or drift; processes with these parameter values typically yield longer first-passage times. Conversely, patients experiencing early failure are assigned higher $D$ or $\mu$ and initial conditions closer to the barrier, consistent with shorter first-passage times.
@@
Final remarks.  We conclude this section by noting a phenomenon visible in Figs. 3,  4: the apparent collinearity between process parameters (or their logarithms). This effect, observed in both Lévy and inverse Gaussian models, leads to data points concentrating along one-dimensional manifolds in parameter space. As reported in the FHT literature [12], this behavior can be attributed to the limited complexity of publicly available datasets rather than to intrinsic shortcomings of the models themselves. We reinforce this notion by showing that the parameter space of the synthetic NonPH dataset does not exhibit this effect (see Appendix C).
Figure 4: Feature–parameter relationships in the Lévy and inverse Gaussian DeepFHT models. Top: Framingham dataset with Lévy model, showing the parameter space $\{x_{0},D\}$ colored by systolic (left) and diastolic (right) blood pressure. Bottom: GBSG2 dataset with inverse Gaussian model, showing the parameter space $\{x_{0},\mu\}$ for patients with grade 1 (left) and grade 3 (right) tumors. In both cases, clinical risk factors align with model-derived high-risk regions (small $x_{0}$, large $D$ or $\mu$), supporting the physical interpretability of the parameterization.
@@ Appendix A Model architecture and hyperparameters
Model GBSG2 Framingham SUPPORT2 NonPH
Lévy $0.694\pm 0.017$ $0.705\pm 0.012$ $0.854\pm 0.004$ $0.840\pm 0.004$
Inverse Gaussian $0.692\pm 0.024$ $0.700\pm 0.013$ $0.854\pm 0.005$ $0.837\pm 0.010$
@@
Appendix B Synthetic dataset generation
@@
Appendix C Physics-based explainability: time interpolation plots
Figure 5: Time interpolation in parameter space across models for GBSG2, SUPPORT2 and NonPH datasets. Notice the absence of collinearity between parameters for the synthetic NonPH datasets, as discussed in Sec. V.2